# Routing Guidance for Emerging Transportation Systems with Improved Dynamic Trip Equity

### v2

@@ IV Routing Guidance System Design with Improved DTE
This section proposes a D-RGS that optimally recommends vehicle routes to enhance $DTE_{i}(t)$ while considering potential traffic congestion. Let $d^{i}$ denote the departure time of vehicle $i\!\in\!\mathcal{N}(t)$ from its origin, i.e., $a_{o_{i}}^{i}\!\!=\!d^{i}$. For any vehicle $i$ arriving at a DMP $v_{k}$ at time $t\!=\!a_{k}^{i}$, the D-RGS provides an optimal route recommendation to the vehicle by addressing
@@
The problem (21) subject to (22) can be addressed by dynamic programming [16, 17] each time a vehicle reaches a DMP, and the optimal solution is denoted as $r_{k,d_{i}}^{*}$. It is worth noting that, if vehicle $i$ has no RRCs at $v_{k}$, i.e., $|\mathcal{C}_{i}(a_{k}^{i})|\!=\!1$, the D-RGS then provides the shortest route for vehicle $i$, as $DTE_{i}(a_{k}^{i})\!=\!1$ naturally holds in this case. Unlike one-time optimization, the D-RGS continuously re-optimizes and updates the route recommendation at each MDP by maximizing the dynamic trip equity $DTE_{i}(a_{k}^{i})$. This facilicates incorporating real-time traffic data and coping with environmental uncertainties.
@@
The D-RGS provides socially optimal route recommendations to individual vehicles that enhance trip equity across travelers. In practice, vehicles may deviate from the recommended routes based on their own route optimization metric or preference [19], such as minimizing travel time or selecting routes with better road conditions. Our follow-up work [20] focuses on effective incentive design to improve route compliance and social optimality.
@@ V Simulation Studies
This section performs simulation studies to illustrate the effectiveness of the proposed framework and the D-RGS.
V-A Setup
Figure 1: An urban road network in Boston.
As shown in Fig. 1, we consider an urban road network in Boston that comprises $58$ nodes, including $45$ road intersections (in blue), $8$ origins (in green), and $5$ destinations (in red). The simulation includes $1,000$ vehicles with random departure times between $08\!\!:\!\!00$-$10\!\!:\!\!00$. The OD pair of each vehicle is randomly selected from the origin and destination sets. Based on indices, vehicles are categorized as: private vehicles ($1\!-\!500$), autonomous vehicles ($501\!-\!800$), and ride-hailing vehicles ($801\!-\!1000$), where each ride-hailing vehicle is shared by two travelers, i.e., $m\!=\!2$ in (17a). All vehicles travel at a constant free-flow speed of $27$ km/h, and the nominal travel time on each edge is obtained from OpenStreetMap [21]. The maximum traffic flow on each edge is considered as $5$ vehicles per minute per lane. The parameter settings for each vehicle type satisfying the conditions in Proposition 1 are provided in Table I. To enhance the computational efficiency, Yen’s algorithm is employed to select $7$ shortest routes from a DMP $v_{k}$ to the destination $d_{i}$ when forming the feasible route set $\mathcal{R}_{k,d_{i}}$.
@@ V-A Setup
V-B Results and Analysis
We compare the proposed method with the pre-planned and dynamic shortest-route strategies, where the pre-planned shortest-route strategy minimizes the trip time based on the nominal travel time on each edge, while the dynamic strategy minimizes the travel time dynamically using the monitored and estimated traffic flow information. Traffic congestion is monitored in real-time at intervals of $\Delta{t}\!=\!60$ seconds. The DTX of each vehicle integrating the actual travel time in (10) in the three strategies is shown in Fig. 2. As shown in Fig. 2(a), the pre-planned strategy results in a wide distribution of trips with very high and low dynamic trip indices, while trips in the mid-range are relatively sparse. The dynamic shortest-route strategy (see Fig. 2(b)), which incorporates real-time traffic information and improves performance for some vehicles with higher DTX, yet a significant fraction of vehicles still experiences a small DTX value. In contrast, the proposed method leads to a more balanced distribution, as shown in Fig. 2(c), where trip indices converge closer to the average level. This improvement enhances the DTE for all trips from $0.734$ to $0.818$, demonstrating the effectiveness of our approach. Notably, the DTE here is computed by
$DTE=1-\frac{\sum_{i=1}^{S}\sum_{i^{\prime}=1}^{S}\!|DTX_{i}\!-\!DTX_{i^{\prime}}|}{2S^{2}DTX_{\text{mean}}}.$
Similar to (16), $S\!=\!1200$ denotes the total number of traveler trips. Note that each ride-hailing vehicle serves two travelers. Thus, $DTX_{i+200}\!=\!DTX_{i}$ for $i\!=\!801,\dots,1000$.
@@ V-B Results and Analysis
Figure 4: Comparison of the average trip cost.
The comparison of the average trip time and cost in the three methods are shown in Fig. 3 and Fig. 4, respectively. We observe that, in comparison with the pre-planned shortest-route strategy, our method significantly reduces the average trip time for private, ride-hailing, and autonomous vehicles by approximately $30\%$, $33\%$, and $42\%$, respectively, leading to savings of about \$$\text{\$}5.7$, \$$\text{\$}2.7$, and \$$\text{\$}3.6$ per trip for each type of vehicle. It is also interesting to see that, compared to the dynamic shortest-route strategy, our method slightly increases the average trip time for private vehicles by $3.1$ minutes, leading to \$$\text{\$}0.84$ increase in the average trip cost. However, it contributes to reducing the trip time and cost for both ride-hailing and autonomous vehicles by approximately $8.8\%$ and $14.7\%$, respectively. This aligns with real-world expectations, as private vehicle owners typically belong to higher-income groups, whereas individuals with lower incomes are more likely to take cost-effective travel options such as ride-hailing and autonomous vehicles. The simulation results above show that our method redistributes travel costs across different vehicle types through route optimization, contributing to a more equitable and efficient transportation system.
@@ VI Conclusion
In this paper, we presented a framework to quantify trip quality and equity in a dynamic emerging transportation system, integrating trip time, cost, and convenience. We developed a routing guidance system to optimize vehicle route recommendations, incorporating real-time and anticipated traffic congestion while improving trip equity. In addition, we established conditions that ensured perfect trip equity in free-flow networks. Our simulation studies on a Boston road network demonstrated that the proposed approach increased trip equity from $0.734$ to $0.818$ compared to the pre-planned shortest-route strategy, leading to $11.4\%$ improvement and a more equitable transportation system. Future work will focus on learning compliance probability and designing effective incentive mechanisms to enhance drivers’ alignment with routing recommendations, thereby improving social equity and system-wide optimality.

### v3

@@ IV Routing Guidance System Design with Improved DTE
This section proposes a D-RGS that optimally recommends vehicle routes to enhance $DTE_{i}(t)$. Let $a_{o_{i}}^{i}$ be the departure time of vehicle $i\!\in\!\mathcal{N}(t)$ from its origin. For vehicle $i$ arriving at a DMP $v_{k}$ at time $t\!=\!a_{k}^{i}$, the D-RGS provides an optimal route recommendation to the vehicle by addressing
@@
The problem (21) subject to (22) can be addressed by dynamic programming [15, 16] each time a vehicle reaches a DMP, and the optimal solution is denoted as $r_{k,d_{i}}^{*}$. Note that, if vehicle $i$ has no RRCs at $v_{k}$, i.e., $|\mathcal{C}_{i}(a_{k}^{i})|\!=\!1$, the D-RGS provides the shortest route for it, as $DTE_{i}(a_{k}^{i})\!=\!1$ naturally holds in this case. Unlike one-time optimization, the D-RGS continuously optimizes and updates the route recommendation at each MDP by maximizing the dynamic trip equity $DTE_{i}(a_{k}^{i})$. This facilitates incorporating real-time traffic data and coping with environmental uncertainties.
@@
The D-RGS provides socially optimal route recommendations to individual vehicles that enhance trip equity across travelers. In practice, vehicles may deviate from the recommended routes based on their own optimization metric or preference [18], such as minimizing travel time or selecting routes with better road conditions. Our follow-up work [19] focuses on effective incentive design to improve route compliance and social optimality.
@@ V Simulation Studies
We consider an urban road network in Boston comprising $45$ road intersections, $8$ origin nodes, and $5$ destination nodes. The simulation involves $1,000$ vehicles with random departure times between 08:0008\!\!:\!\!00-10:0010\!\!:\!\!$0$, including $500$ private vehicles, $300$ autonomous vehicles, and $200$ ride-hailing vehicles, where each ride-hailing vehicle serves two travelers (i.e., m=2m\!=\!2 in (17a)). The OD pairs are randomly selected from the origin and destination sets. We assume all vehicles travel at a free-flow speed of $27$ km/h, and the nominal travel time on each edge is obtained from OpenStreetMap [20]. The maximum traffic flow is set as $5$ vehicles per minute per lane. The parameter settings for each vehicle type, satisfying the conditions in Proposition 1, are provided in Table I. To enhance computational efficiency, Yen’s algorithm is employed to generate $7$ shortest routes from a DMP $v_{k}$ to the destination $d_{i}$ when forming the feasible route set $\mathcal{R}_{k,d_{i}}$.
@@
V-A Results and Analysis
We compare the proposed method with the pre-planned and dynamic shortest-route strategies, where the pre-planned shortest-route (P-SR) strategy minimizes the trip time based on the nominal travel time, while the dynamic shortest-route (D-SR) strategy incorporates the monitored and estimated traffic flow information to optimize the trip time dynamically. Traffic congestion is monitored in real-time at intervals of $\Delta{t}\!=\!60$ seconds. The DTX of each vehicle, integrating the actual travel time in (10) under the three strategies, is shown in Fig. 1. The DTEs for all completed trips in P-SR, D-SR, and our methods are $0.734$, $0.777$, and $0.818$, respectively. As we observe, the P-SR strategy results in a wide spread of DTX values, indicating large inequity across individual trips. The D-SR strategy improves performance by increasing the median DTX, but still exhibits considerable variability, with a substantial number of vehicles experiencing low trip equity. In contrast, our proposed method achieves a more centralized distribution of DTX across all vehicle types, showing a more consistent and balanced trip equity.
@@ V-A Results and Analysis
Fig. 2 compares the average trip time in the three strategies. As is shown, the proposed method significantly reduces the average trip time for private, ride-hailing, and autonomous vehicles by approximately $0\%, 33$$3\%$, and $42\%$, respectively, leading to per-trip savings of about $\text{\$}5.7$, $\text{\$}2.7$, and $\text{\$}3.6$ for each vehicle type. Compared to the D-SR strategy, our method increases the average trip time for private vehicles by $3.1$ minutes, while reducing trip times for both ride-hailing and autonomous vehicles by approximately $8.8\%$ and $14.7\%$, respectively. This outcome aligns with real-world expectations, where private vehicle owners typically belong to higher-income groups, and a slight increase in their travel time as well as travel costs leads to equity enhancement for other travelers. These results demonstrate that our method effectively redistributes travel costs across vehicle types through route optimization, contributing to a more equitable and efficient transportation system.
@@ VI Conclusion
In this paper, we presented a framework to quantify trip quality and equity in a dynamic emerging transportation system, integrating trip time, cost, and convenience. We developed a routing guidance system to optimize vehicle route recommendations, incorporating real-time and anticipated traffic congestion while improving trip equity. In addition, we established conditions that ensured perfect trip equity in free-flow networks. Our simulation studies on a Boston road network demonstrated that the proposed approach increased trip equity from $0.734$ to $0.818$ compared to the pre-planned shortest-route strategy, leading to $11.4\%$ improvement and a more equitable transportation system. Future work will focus on learning the compliance rate and designing effective incentive mechanisms to enhance drivers’ alignment with routing recommendations.